# MUDAS: Mote-scale Unsupervised Domain Adaptation in Multi-label Sound Classification

### v1

@@ V-B Selection for Data Used in Relearning
  • Limited gains from increasing target dataset size To assess the impact of target dataset size on accuracy, we conducted experiments with a fixed source dataset size of 500 and varied target dataset sizes: 500 (1:1), 1000 (2:1), 1500 (3:1), 2000 (4:1), 2500 (5:1), and 3000 (6:1). The analysis was performed using two distinct target locations and a single source location. As illustrated on the left side of Figure 3, increasing the target dataset size led to only marginal improvements in accuracy, with Micro AUPRC fluctuations of about 0.07 in Brooklyn and 0.02 in Lower Manhattan. Specifically, Brooklyn initially saw a decline in accuracy, followed by a gradual increase as the target-to-source ratio shifted from 3:1 to 6:1. In contrast, Lower Manhattan showed a steady decline in accuracy as the target dataset size grew, but the variation was less pronounced than in Brooklyn. This suggests that for a given source location, a modest target-to-source ratio is sufficient for effective adaptation. In our evaluations, a ratio of 1:1 or 2:1 was found to be effective.
@@
  • Impact of probability range In real-world scenarios, label information may not always be available for the target dataset. Therefore, we explored the use of probability scores as an alternative source of information, under the assumption that higher probabilities are more likely to correspond to correctly labeled data. Our MUDAS framework leverages probability scores from classifier inference. We trained the model using only the source dataset and used it to infer probability scores for the target dataset. These probabilities were then normalized by dividing each class’s score by the range between its maximum and minimum values, creating three sets based on average probability: high, medium, and low. The 500 target samples were categorized accordingly. We tested MUDAS on these three target sets, and as shown in Table I, the high-probability set consistently achieved significantly higher Micro AUPRC scores compared to the medium and low-probability sets. These results demonstrate that data selection based on high-probability target samples leads to better performance than selecting from medium or low-probability samples.
@@
$D=\frac{1}{k}\sum_{i=1}^{k}(\frac{w_{i}\cdot z^{wt}_{i,j}}{\max_{j=1}^{m}z^{ws}_{i}-\min_{j=1}^{m}z^{ws}_{i}})$ (14)
@@
Once the normalization is applied, the target probability $z^{wt}_{i,j}$ is multiplied by the class weight $w_{i}=\frac{\text{Number of labels in class $i$}}{\text{Total number of all class labels}}$. This weighting assigns higher importance to classes with more frequent labels in the source dataset. The weighted values for each class are then combined and averaged to obtain the value $D$. Finally, data selection is completed by ranking the values in $D$ from highest to lowest.
@@
VI EVALUATION
In this section, we evaluate the performance of MUDAS. We use Version 2 of the SONYC-UST dataset [7]. This dataset provides latitude and longitude coordinates for each recording, allowing us to organize samples by specific locations across New York City. It includes 8 coarse-grained and 23 fine-grained sound classes; for this study, we focus on the broader coarse-grained classes to simplify the analysis and improve generalizability.
@@ VI EVALUATION
To enhance feature diversity, we apply both strong and weak data augmentations. Strong augmentation is achieved using SpecAugment [30], which generates meaningful variations in the audio spectrograms, while weak augmentation consists of simple x-axis flip-flop transformations. Feature embeddings are generated using SONYC-L3 [14], a lightweight embedding model optimized for deployment on mote-scale devices, which ensures efficient processing for resource-constrained environments.
@@
We evaluate MUDAS’s robustness across several scenarios: varying target locations, hyperparameter tuning (exploring different values for the positive threshold $\tau^{+}$ and negative threshold $\tau^{-}$, batch size), and conducting ablation studies to investigate the contributions of individual components to overall performance. Furthermore, we compare MUDAS with AdaMatch [5], a framework designed for multi-class classification. Although this comparison is not an apples-to-apples comparison as the models are trained differently depending on the multi-label or the multi-class objective, it establishes two distinct baseline configurations: one for AdaMatch in multi-class classification settings and another for MUDAS in multi-label classification settings. This enables us to evaluate the relative improvements achieved by MUDAS, despite the differences in problem settings.
@@
TABLE II: Performance of MUDAS across locations (6-class evaluation)
Method Metric Location
@@
Lower-Bound for MUDAS Micro-AUPRC 0.657 0.733 0.479 0.551
@@
Upper-Bound for MUDAS Micro-AUPRC 0.778
@@
Location-based Evaluation. To evaluate MUDAS’s performance, we selected four distinct locations as source domains, each contributing 500 labeled samples for training, with Washington Square as the target domain providing 500 unlabeled samples. We set $\tau^{+}=0.9$ and $\tau^{-}=0.9$ with a batch size of 64. MUDAS’s performance was compared against lower-bound and upper-bound baselines, as shown in Table II.
@@
Figure 4: Impact of varying user-defined parameters for positive and negative thresholds, $\tau^{+}$ and $\tau^{-}$ (6-class evaluation)
TABLE III: Ablation Study: impact of negative threshold and diversity regularization in MUDAS Performance (Micro AUPRC) (6-classes evaluation)
MUDAS w/ Positive Threshold Negative Threshold Diversity Regularization Location
@@
O X X 0.686$\pm$.020 0.735$\pm$.005 0.482$\pm$.012 0.582$\pm$.003
O O X 0.691$\pm$.019 0.736$\pm$.006 0.492$\pm$.015 0.582$\pm$.004
O X O 0.693$\pm$.019 0.739$\pm$.005 0.493$\pm$.012 0.572$\pm$.004
O O O 0.695$\pm$.018 0.740$\pm$.007 0.499$\pm$.013 0.570$\pm$.006
Figure 5: Ablation study: Impact of data selection on MUDAS performance (6-class evaluation)
TABLE IV: Performance comparison between AdaMatch and the baseline in multi-label and multi-class classification settings (3-class evaluation)
Union Square Central Park Brooklyn Lower Manhattan
Baseline for
Multi-Class
AdaMatch
Baseline for
Multi-Label
MUDAS
Baseline for
Multi-Class
AdaMatch
Baseline for
Multi-Label
MUDAS
Baseline for
Multi-Class
AdaMatch
Baseline for
Multi-Label
MUDAS
Baseline for
Multi-Class
AdaMatch
Baseline for
Multi-Label
MUDAS
Metric
Micro-AUPRC 0.732$\pm$.007 0.749$\pm$.008 0.763$\pm$.014 0.801$\pm$.005 0.738$\pm$.007 0.741$\pm$.004 0.824$\pm$.007 0.826$\pm$.005 0.646$\pm$.006 0.666$\pm$.008 0.575$\pm$.020 0.642$\pm$.010 0.661$\pm$.009 0.675$\pm$.008 0.665$\pm$.006 0.668$\pm$.007
Class Label
Engine 0.710$\pm$.006 0.725$\pm$.006 0.744$\pm$.018 0.790$\pm$.020 0.716$\pm$.006 0.718$\pm$.004 0.856$\pm$.008 0.851$\pm$.010 0.651$\pm$.004 0.661$\pm$.007 0.627$\pm$.012 0.681$\pm$.007 0.662$\pm$.004 0.667$\pm$.006 0.693$\pm$.013 0.682$\pm$.005
Alert signal 0.776$\pm$.016 0.774$\pm$.010 0.749$\pm$.012 0.749$\pm$.019 0.758$\pm$.012 0.760$\pm$.011 0.756$\pm$.009 0.772$\pm$.012 0.339$\pm$.013 0.353$\pm$.020 0.280$\pm$.005 0.267$\pm$.019 0.705$\pm$.013 0.703$\pm$.006 0.651$\pm$.041 0.714$\pm$.010
Human voice 0.838$\pm$.003 0.847$\pm$.006 0.856$\pm$.012 0.875$\pm$.017 0.877$\pm$.006 0.876$\pm$.006 0.873$\pm$.005 0.864$\pm$.005 0.826$\pm$.015 0.853$\pm$.010 0.679$\pm$.030 0.768$\pm$.002 0.853$\pm$.017 0.855$\pm$.013 0.839$\pm$.021 0.849$\pm$.017
Ablation Study. In the ablation study, we systematically analyzed the effects of incorporating negative thresholds, diversity regularization, and data selection by comparing model performance across different configurations. Table III presents the results for MUDAS, beginning with the baseline configuration that uses only a positive threshold and incrementally adding negative thresholds and diversity regularization.
@@
To assess MUDAS’s and AdaMatch’s adaptability in a multi-label context, we establish two separate baseline configurations: one for multi-class classification (AdaMatch) and one for multi-label classification (MUDAS). To compare the performance of each method, two baseline approaches are employed. The first baseline operates in a multi-class classification setting, utilizing a softmax activation function and categorical cross-entropy as the loss function. The second baseline is designed for a multi-label classification setting, employing a sigmoid activation function and binary cross-entropy for the loss function. Both baselines are trained exclusively on source-labeled data, providing a benchmark for evaluating the proposed methods.
In this evaluation, we set $\tau=0.9$ for AdaMatch and $\tau^{+}=0.9$, $\tau^{-}=0.9$ for MUDAS, maintaining a consistent batch size of 64 for both models. To address measurement bias due class imbalance—where multi-class classification often does not predict underrepresented classes, leading to artificially inflated performance—we select three classes with a higher number of labels (engine, alert signal, and human voice) to reduce this bias.
We evaluated the performance of the models using data from four different locations, with the results presented in Table IV. The evaluation shows that AdaMatch consistently outperforms its baseline for multi-class classification, while MUDAS consistently outperforms its baseline for multi-label classification. These results highlight that both AdaMatch and MUDAS are effective unsupervised domain transfer techniques, surpassing their respective baselines.
MUDAS tends to outperform AdaMatch, but is not always superior. More specifically, in Table IV, in Union Square and Central Park, we observe that MUDAS significantly outperforms AdaMatch in Micro-AUPRC, with improvements of 0.052 and 0.085, respectively. In Lower Manhattan, MUDAS and AdaMatch exhibit nearly identical performance, with their AUPRC difference falling within the margin of error. But, in Brooklyn, MUDAS slightly underperforms compared to AdaMatch, with a difference of 0.024.
Interestingly, a similar trend is observed with the baselines. In Union Square and Central Park, the multi-label baseline achieves higher Micro-AUPRCs compared to the multi-class baseline. In Lower Manhattan, the AUPRC difference between the two baselines is within the error range. But, in Brooklyn, the multi-class baseline performs better.
@@
VII CONCLUSIONS
@@
Extensive evaluations using the SONYC-UST dataset validate MUDAS’s ability to handle domain shifts effectively, achieving superior performance compared to baseline models for multi-label urban sound classification. Importantly, MUDAS achieves this while adhering to the practical limitations of real-world IoT environments. These results underscore MUDAS’s potential to scale and adapt UDA methods for complex and dynamic settings. Note that our evaluation of MUDAS assumes training from scratch. However, in practical scenarios, as described in Section V-A, MUDAS ought to be initialized with a pre-trained source model, which would further reduce the reliance on extensive source datasets and save valuable storage space on resource-constrained devices.

### v2

@@ 5.2. Selection for Data Used in Relearning
  • Adequate adaptation gains from modest target dataset size To assess the impact of target dataset size on accuracy, we conducted experiments with a fixed source dataset of 500 and varied target datasets: 500 (1:1), 1000 (2:1), 1500 (3:1), 2000 (4:1), 2500 (5:1), and 3000 (6:1), using two target locations and one source location. As shown in Figure 3 (left), increasing the target dataset size led to modest accuracy changes: Micro AUPRC fluctuated by 0.07 in Brooklyn and 0.02 in Lower Manhattan. These results indicate that a modest target-to-source ratio, such as 1:1 or 2:1, is sufficient for effective adaptation and minimize performance degradation.
@@
  • Impact of probability range In real-world scenarios, label information may not always be available for the target dataset. Therefore, we explored the use of probability scores as an alternative source of information, under the assumption that higher probabilities are more likely to correspond to correctly labeled data. Our MUDAS framework leverages probability scores from classifier inference. We trained the model using only the source dataset and used it to infer probability scores for the target dataset. These probabilities were then normalized by dividing each class’s score by the range between its maximum and minimum values, creating three sets based on average probability: high, medium, and low. The 500 target samples were categorized accordingly. We tested MUDAS on these three target sets, and as shown in Table 1, the high-probability set consistently achieved significantly higher Micro AUPRC scores compared to the medium and low-probability sets. These results demonstrate that data selection based on high-probability target samples leads to better performance than selecting from medium or low-probability samples.
@@
(14) $D=\frac{1}{k}\sum_{i=1}^{k}(\frac{w_{i}\cdot z^{wt}_{i,j}}{\max_{j=1}^{m}z^{ws}_{i}-\min_{j=1}^{m}z^{ws}_{i}})$
@@
[…]
@@
6. EVALUATION
In this section, we evaluate the performance of MUDAS. We use Version 2 of the SONYC-UST dataset (SONYC_UST_V2). This dataset provides latitude and longitude coordinates for each recording, allowing us to organize samples by specific locations across New York City. It includes 8 coarse-grained and 23 fine-grained sound classes; for this study, we focus on the broader coarse-grained classes to simplify the analysis and improve generalizability.
@@ 6. EVALUATION
To enhance feature diversity, we apply both strong and weak data augmentations. Strong augmentation is achieved using SpecAugment (SpecAugment), which generates meaningful variations in the audio spectrograms, while weak augmentation consists of simple x-axis flip-flop transformations. Feature embeddings are generated using SONYC-L3 (MKII), a lightweight embedding model optimized for deployment on mote-scale devices, which ensures efficient processing for resource-constrained environments.
@@
We evaluate MUDAS’s robustness across several scenarios: varying target locations, hyperparameter tuning (exploring different values for the positive threshold $\tau^{+}$ and negative threshold $\tau^{-}$, batch size), and conducting ablation studies to investigate the contributions of individual components to overall performance. Furthermore, we compare MUDAS with AdaMatch (AdaMatch), a framework designed for multi-class classification. Although this comparison is not an apples-to-apples comparison as the models are trained differently depending on the multi-label or the multi-class objective, it establishes two distinct baseline configurations: one for AdaMatch in multi-class classification settings and another for MUDAS in multi-label classification settings. This enables us to evaluate the relative improvements achieved by MUDAS, despite the differences in problem settings.
@@
Table 2. Performance of MUDAS in various source locations (6-class evaluation)
Method Metric Source Location
@@
Lower-Bound
for MUDAS
Micro-AUPRC 0.657 0.733 0.479 0.551
@@
Upper-Bound
for MUDAS
Micro-AUPRC 0.778
@@
Figure 4. Impact of varying user-defined parameters for positive and negative thresholds, $\tau^{+}$ and $\tau^{-}$ (6-class evaluation)
Location-based Evaluation. To evaluate MUDAS’s performance, we selected four distinct locations as source domains, each contributing 500 labeled samples for training, with Washington Square as the target domain providing 500 unlabeled samples. We set $\tau^{+}=0.9$ and $\tau^{-}=0.9$ with a batch size of 64. MUDAS’s performance was compared against lower-bound and upper-bound baselines, as shown in Table 2.
@@
Table 3. Ablation Study: impact of negative threshold and diversity regularization in MUDAS Performance (Micro AUPRC) (6-classes evaluation)
MUDAS w/
Positive
Threshold
Negative
Threshold
Diversity
Regularization
Source Location
@@
[…]
Figure 5. Ablation study: Impact of data selection on MUDAS performance (6-class evaluation)
Ablation Study. In the ablation study, we systematically analyzed the effects of incorporating negative thresholds, diversity regularization, and data selection by comparing model performance across different configurations. Table 3 presents the results for MUDAS, beginning with the baseline configuration that uses only a positive threshold and incrementally adding negative thresholds and diversity regularization.
@@
Table 4. Performance comparison between AdaMatch and the baseline in multi-class classification settings (3-class evaluation)
Source Location
Union
Square
Central
Park
Brooklyn
Lower
Manhattan
Metric
[…]
Class Label
[…]
@@
To assess MUDAS’s and AdaMatch’s adaptability in a multi-label context, we establish two separate baseline configurations: one for multi-class classification (AdaMatch) and one for multi-label classification (MUDAS). To evaluate the performance of each method in isolation, two different baseline approaches are employed. The first baseline operates in a multi-class classification setting, utilizing a softmax activation function and categorical cross-entropy as the loss function. The second baseline is designed for a multi-label classification setting, employing a sigmoid activation function and binary cross-entropy for the loss function. Both baselines are trained exclusively on source-labeled data, providing a benchmark for evaluating the proposed methods.
In this evaluation, we set $\tau=0.9$ for AdaMatch and $\tau^{+}=0.9$, $\tau^{-}=0.9$ for MUDAS, maintaining a consistent batch size of 64 for both models. To reduce measurement bias due to class imbalance—wherein multi-class classification often does not predict underrepresented classes, leading to artificially inflated performance—we limit the set of classes to those that have a sufficient number of labels. Accordingly, 3 classes are selected for a fairer comparative evaluation, namely engine, alert signal, and human voice.
We evaluated the performance of the models using data from four different locations, with the results presented in Table 4, 5. The evaluation shows that AdaMatch consistently outperforms its baseline for multi-class classification, while MUDAS consistently outperforms its baseline for multi-label classification. Relatively speaking, AdaMatch shows improvements of up to 0.020, whereas MUDAS shows stronger improvements of up to 0.067. These results corroborate that both AdaMatch and MUDAS are effective unsupervised domain transfer techniques.
@@
Table 5. Performance comparison between MUDAS and the baseline in multi-label classification settings (3-class evaluation)
Source Location
Union
Square
Central
Park
Brooklyn
Lower
Manhattan
Metric
[…]
Class Label
[…]
Feasibility of Deployment on Edge Devices Although MUDAS has been evaluated in an emulated environment, the derived model can be readily deployed on existing edge devices. Each 1-second embedding, which would be sufficient for accurate noise classification, needs only 1 KB. Thus, when using 500 source and target embeddings, the total memory requirement is just over 2 MB of flash storage. Moreover, the model utilizes only 141 KB of activation memory and 67 KB of parameter memory during training, highlighting its suitability for deployment on resource-constrained edge platforms.
7. CONCLUSIONS
@@
Extensive evaluations using the SONYC-UST dataset validate MUDAS’s ability to handle domain shifts effectively, achieving superior performance compared to baseline models for multi-label urban sound classification. Importantly, MUDAS achieves this while adhering to the practical limitations of real-world IoT environments. These results underscore MUDAS’s potential to scale and adapt UDA methods for complex and dynamic settings. Note that our evaluation of MUDAS assumes training from scratch. However, in practical scenarios, as described in Section 5.1, MUDAS ought to be initialized with a pre-trained source model, which would further reduce the reliance on extensive source datasets and save valuable storage space on resource-constrained devices.